OpenStudy (anonymous):

Simplify: Square root of 8xy^2

OpenStudy (tennistar):

Um... that number? you need numbers and not letters or symbols

OpenStudy (anonymous):

what?!

OpenStudy (anonymous):

√8xy^2

OpenStudy (anonymous):

that's my question ^^^^ ha, sorry!

OpenStudy (anonymous):

\[(8xy ^{2})^{\frac{ 1 }{ 2 }}\] The rules of indices tells us: \[(x^a)^b=x ^{ab}\] Can you go from there?

OpenStudy (anonymous):

and \[(xyz)^a=x^ay^az^a\]

OpenStudy (anonymous):

um... I'm not sure